• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 06:06:59 on May19,2008

    Entry number 238723

    Followups:

    keyword=Target: brief alarms

    Target alarmed - the ESR supply temperature (haCryo_Supply_T)
    went above the yellow and red alarm limits (both are set to 14.50 K
    at present); it went up to 14.54 K. High power heater was down to zero W. Tried to open up JT valve two steps of 1%. Cell had
    warmed by 0.2 K to 22.2 K. Cycled HPH from PID to manual (10 W) back to PID. After about 2 minutes, the 15 K supply temperature went
    back down below the alarm point (now 13.95 K), and cell cooled
    back down to 20.00K, and HPH turned back on.

    After another minute or so, returned JT valve to original
    setpoint, and all is back to normal.